OpenStudy (anonymous):

Solve. -1/4(x-8)+1/2(x+2)=x-9

OpenStudy (mertsj):

Multiply both sides by 4 to clear out the fractions.

OpenStudy (mertsj):

\[4[−\frac{1}{4}(x−8)]+4[\frac{1}{2}(x+2)]=4(x−9)\]

OpenStudy (mertsj):

\[-1(x-8)+2(x+2)=4x-36\]

OpenStudy (mertsj):

Finish it.

OpenStudy (anonymous):

4x/4-36/4 x=8

OpenStudy (mertsj):

no